OpenStudy (kaz2019):

Please help I am losing the will to live!! Raw scores on behavioral tests are often transformed for easier comparison. A test of reading ability has mean 51 and standard deviation 10 when given to third graders. Sixth graders have mean score 83 and standard deviation 11 on the same test. To provide separate "norms" for each grade, we want scores in each grade to have mean 100 and standard deviation 20. (Round your answers to two decimal places.) a) What linear transformation will change third-grade scores x into new scores Xnew = a + bx that have the desired mean and standard deviation?

jimthompson5910 (jim_thompson5910):

Focus on the third grade class. We have mean = 51 standard deviation = 10

jimthompson5910 (jim_thompson5910):

we want the following mean = 100 standard deviation = 20

jimthompson5910 (jim_thompson5910):

so we want the 51 to turn into 100 we also want the 10 to turn into 20 we essentially have 2 points on a line: (51,100) and (10,20)

jimthompson5910 (jim_thompson5910):

are you able to find the equation of the line through those two points?

OpenStudy (kaz2019):

I cant remember the formula for that but I thought we had to use the Xnew=a+bx?

jimthompson5910 (jim_thompson5910):

do you know how to find the slope of a line using the slope formula? does this formula look familiar? Slope Formula \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (kaz2019):

Yes I remember how to do that. \[m=\frac{ 20-100 }{ 10-51 }=\frac{ -80 }{ -41 }=\frac{ 80 }{ 41 }\]

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now let's pick one of the two points, let's say (10,20) this means x = 10 and y = 20

jimthompson5910 (jim_thompson5910):

we'll plug these three values m = 80/41 x = 10 y = 20 into this slope intercept formula y = mx+b then solve for b

jimthompson5910 (jim_thompson5910):

\[\Large y = mx+b\] \[\Large 20 = \frac{80}{41}*10+b\] solve for b

OpenStudy (kaz2019):

That would come out to \[\frac{ 20 }{ 41 }\]

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

m = 80/41 b = 20/41 so the equation y = mx+b turns into \[\Large y = \frac{80}{41}x+\frac{20}{41}\]

jimthompson5910 (jim_thompson5910):

Therefore, \[\Large X_{\text{new}} = a+b*x\] turns into \[\Large X_{\text{new}} = \frac{20}{41}+\frac{80}{41}*x\]

OpenStudy (kaz2019):

Im trying to understand everything and i feel really confused on how we went from slope formula to this

jimthompson5910 (jim_thompson5910):

well y = mx+b and a+bx are the same idea the 'm' in the first equation and the 'b' in the second equation are the slope. It's unfortunate that 'b' is used in two different ways here. I can see how it's confusing

jimthompson5910 (jim_thompson5910):